What is the solve for unknwon: log^3 (2X + 1) = 1
Answer by lwsshak3(6513)

log^3 (2X + 1) = 1
convert to exponential form: Base(3) raised to log of number(1)=number(2x+1)
3^1=2x+1=3
2x=2
x=1